ZYZ Inc. is considering a project with the following cash flows:

Year

Cash Flow (CF)

0

-$200,000

1

$30,000

2

$40,000

3

$50,000

4

$60,000

5

$70,000

If the discount rate is 5%, what is the NPV of the proposed project?

Question 3 options:

$11,572.99

$12,253.47

$21,009.43

$10,572.99
Answer #1

Year a Cash Flow PV Factor PV Of Cash Flow b c=1/1.05^a d=b*c o $ -2,00,000 1 $ -2,00,000.00 1 $ 30,000 0.95238095 $ 28,571.4